PFIFOFAST(8) Linux PFIFOFAST(8) NAME pfifofast−three-bandfirstin,firstoutqueue DESCRIPTION pfifofastisthedefaultqdiscofeachinterface. Wheneveraninterfaceiscreated,thepfifofastqdiscisautomaticallyusedasaqueue.Ifanother qdiscisattached,itpreemptsthedefaultpfifofast,whichautomaticallyreturnstofunctionwhen anexistingqdiscisdetached. Inthissensethisqdiscismagic,andunlikeotherqdiscs. ALGORITHM Thealgorithmisverysimilartothatoftheclassfultc-prio(8)qdisc.pfifofastislikethreetc-pfifo(8)queuessidebyside,wherepacketscanbeenqueuedinanyofthethreebandsbasedon theirTypeofServicebitsorassignedpriority. Notallthreebandsaredequeuedsimultaneously-aslongaslowerbandshavetraffic,higher bandsareneverdequeued.Thiscanbeusedtoprioritizeinteractivetrafficorpenalize’lowest cost’traffic. Eachbandcanbetxqueuelenpacketslong,asconfiguredwithifconfig(8)orip(8).Additional packetscominginarenotenqueuedbutareinsteaddropped. Seetc-prio(8)forcompletedetailsonhowTOSbitsaretranslatedintobands. PARAMETERS txqueuelen Thelengthofthethreebandsdependsontheinterfacetxqueuelen,asspecifiedwith ifconfig(8)orip(8). BUGS Doesnotmaintainstatisticsanddoesnotshowupintcqdiscls.Thisisbecauseitistheautomaticdefaultintheabsenceofaconfiguredqdisc. SEEALSO tc(8) AUTHORS AlexeyN.Kuznetsov,<kuznet@ms2.inr.ac.ru> Thismanpagemaintainedbyberthubert<ahu@ds9a.nl> iproute2 10January2002 1